# Covariant Fractons and Weitzenböck Torsion

### v1

@@ 7 BRST formulation
In this section we formulate the brst rules (2.1) defining the covariant fracton gauge theory in an equivalent way, showing in which sense the covariant fracton gauge theory can be deduced from linearised mhs theory.
Introduce two doublets $(B_{\mu\nu},\tilde{b}_{\mu\nu})$ and $(r_{\mu},\tilde{\xi}_{\mu})$, with the following transformations
@@
requiring that both $\tilde{b}_{\mu\nu}$ and $r_{\mu}$ are commuting with ghost number zero, so that $B_{\mu\nu}$ is anticommuting with ghost number $-1$ and $\tilde{\xi}_{\mu}$ is anticommuting with ghost number one. The local $s$-cohomology on the jet space $\{h^{(\text{fr})}_{\mu\nu},\lambda\}$ is equivalent to the local $s$-cohomology on the enlarged space including the two doublets $\{h^{(\text{fr})}_{\mu\nu},\lambda,B_{\mu\nu},\tilde{b}_{\mu\nu},r_{\mu},\tilde{\xi}_{\mu}\}$. If we redefine $\tilde{b}_{\mu\nu}$ and $\tilde{\xi}_{\mu}$ in this way
@@
$s\,b_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}-\partial_{\nu}\,\mathring{\xi}_{\mu},\quad s\,\mathring{\xi}_{\mu}=0,$ (7.3)
@@
The local $s$-cohomology on $\{h_{\mu\nu}^{\text{(fr)}},\lambda\}$ is equivalent to that on $\{B_{\mu\nu},h_{\mu\nu},b_{\mu\nu},r_{\mu},\mathring{\xi}_{\mu},\lambda\}$. In other words, the doublet $(r_{\mu},\tilde{\xi}_{\mu})$ implements the condition “$\mathring{\xi}_{\mu}=\tfrac{1}{2}\,\partial_{\mu}\,\lambda$”, which is implied by “$r_{\mu}=0$”; and the doublet $(B_{\mu\nu},\tilde{b}_{\mu\nu})$ implements the condition “$b_{\mu\nu}=\partial_{\mu}\,r_{\nu}-\partial_{\nu}\,r_{\mu}$”, which is implied by “$B_{\mu\nu}=0$”. In this sense, the covariant fracton gauge transformations can be thought as the restriction of linearised diffeomorphisms to the longitudinal ones, as they were firstly introduced Blasi:2022mbl , and covariant fracton gauge theory can be thought as the restriction of the linearised mhs theory in the case in which the antisymmetric part of the linearised vielbein is $\mathrm{d}$-exact. As a consequence, the linearised torsion is cohomologically equivalent to the fracton field strength in this formulation. Indeed, the linearised vielbein reads
@@
This is the reason why the constants $\alpha_{1}$ and $\alpha_{2}$ appear only in the combination $2\,\alpha_{1}+\alpha_{2}$ n embedding covariant fracton gauge theory in linearised mhs theory.
@@ Acknowledgments
The author is deeply indebted with F. Fecit and C. Imbimbo for encouragements and illuminating discussions and observations.
@@ Appendix A Weitzenböck geometry within Cartan formalism
In the formalism introduced by Cartan, one considers the vielbein one-form $e^{a}=e_{\mu}{}^{a}\,\mathrm{d}x^{\mu}$ and the spin connection one-form $\omega^{a}{}_{b}=\omega_{\mu}{}^{a}{}_{b}\,\mathrm{d}x^{\mu}$, and define the torsion two-form $T^{a}=\frac{1}{2}\,T_{\mu}{}^{a}{}_{\nu}\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}$ and the curvature two-form $R^{a}{}_{b}=\frac{1}{2}\,R_{\mu\nu}{}^{a}{}_{b}\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}$ in the following way:
@@
$R^{a}{}_{b}(\omega)$ $=\mathrm{d}\,\omega^{a}{}_{b}+\omega^{a}{}_{c}\,\omega^{c}{}_{b}.$ (A.1b)
$\omega_{\mu}{}^{a}{}_{b}$ serves as a frame connection (for flat indices). The basis or affine connection $\Gamma_{\mu}{}^{\varrho}{}_{\nu}$ (for curved indices) is related to $\omega_{\mu}{}^{a}{}_{b}$ by the requirement that the full covariant derivative of the vielbein, both with respect to $\omega_{\mu}{}^{a}{}_{b}$ and with respect to $\Gamma_{\mu}{}^{\varrho}{}_{\nu}$ (“second vielbein postulate”), vanishes [^9] This is equivalent to define the covariant derivative with respect to the affine connection $\nabla_{\mu}(\Gamma)$ in terms of the covariant derivative with respect to the spin connection $\nabla_{\mu}(\omega)$ in this way: $\nabla_{\mu}(\Gamma)\,(e^{\varrho}{}_{a}\,V^{a})=e^{\varrho}{}_{a}\,\nabla_{\mu}(\omega)\,V^{a}$, for any $V^{a}$. Indeed on a side $\nabla_{\mu}(\Gamma)\,(e^{\varrho}{}_{a}\,V^{a})=\partial_{\mu}\,V^{\varrho}+\Gamma_{\mu}{}^{\varrho}{}_{\nu}\,V^{\nu}$; on the other $e^{\varrho}{}_{a}\,\nabla_{\mu}(\omega)\,V^{a}=\partial_{\mu}\,V^{\varrho}+e^{\varrho}{}_{a}\,(\partial_{\mu}\,e_{\nu}{}^{a}+\omega_{\mu}{}^{a}{}_{b}\,e_{\nu}{}^{b})\,V^{\nu}$, so that we obtain (A.2). VanNieuwenhuizen:1981ae :
@@
$\ignorespaces T_{\mu}{}^{\varrho}{}_{\nu}(\Gamma)$ $=\Gamma_{[\mu}{}^{\varrho}{}_{\nu]},$ (A.4a)
$R^{\varrho}{}_{\sigma\mu\nu}(\Gamma)$ $=\partial_{[\mu}\,\Gamma_{\nu]}{}^{\varrho}{}_{\sigma}+\Gamma_{[\mu}{}^{\varrho}{}_{\alpha}\,\Gamma_{\nu]}{}^{\alpha}{}_{\sigma},$ (A.4b)
@@
$\ignorespaces T_{\mu}{}^{\varrho}{}_{\nu}(\Gamma(e,\omega))$ $=T_{\mu}{}^{a}{}_{\nu}(e,\omega)\,e^{\varrho}{}_{a},$ (A.5a)
$R^{\varrho}{}_{\sigma\mu\nu}(\Gamma(e,\omega))$ $=R_{\mu\nu}{}^{a}{}_{b}(\omega)\,e^{\varrho}{}_{a}\,e_{\sigma}{}^{b}.$ (A.5b)
@@
$\Gamma_{\mu}{}^{\varrho}{}_{\nu}(e,\omega^{\text{(flat)}})=e^{\varrho}{}_{a}\,\partial_{\mu}\,e_{\nu}{}^{a}+{\omega^{(\text{flat})}}_{\mu}{}^{a}{}_{b}\,e^{\varrho}{}_{a}\,e_{\nu}{}^{b}=(\lambda_{c}{}^{b}\,e^{\varrho}{}_{b})\,\partial_{\mu}\,(\lambda^{c}{}_{a}\,e_{\nu}{}^{a})$ (A.10)
@@
A useful tensor is the contorsion, which is defined as the difference between an arbitrary affine connection and the Levi-Civita connection [^11] Using (A.3), we can also write the contorsion in terms of an arbitrary spin connection and the torsionless one: $-K_{\mu}{}^{\varrho}{}_{\nu}=(\omega_{\mu}{}^{a}{}_{b}-{\omega^{\textsc{(gr)}}}_{\mu}{}^{a}{}_{b})\,e^{\varrho}{}_{a}\,e_{\nu}{}^{b}$.
$-K_{\mu}{}^{\varrho}{}_{\nu}(\Gamma):=\Gamma_{\mu}{}^{\varrho}{}_{\nu}-{\Gamma^{(\textsc{gr})}}_{\mu}{}^{\varrho}{}_{\nu},$ (A.13)
@@
We can relate the Levi-Civita connection, the Weitzenböck connection and the Weitzenböck torsion in the following way. On a side, using (A.13),
$-K_{\mu}{}^{\varrho}{}_{\nu}(W)=W_{\mu}{}^{\varrho}{}_{\nu}-{\Gamma^{\textsc{(gr)}}}_{\mu}{}^{\varrho}{}_{\nu}.$ (A.16)
@@
$-K_{\mu}{}^{\varrho}{}_{\nu}(W)=\tfrac{1}{2}\,(T_{\mu}{}^{\varrho}{}_{\nu}(W)-T_{\mu\nu}{}^{\varrho}(W)+T^{\varrho}{}_{\mu\nu}(W)).$ (A.17)
@@
${\Gamma^{\textsc{(gr)}}}_{\mu}{}^{\varrho}{}_{\nu}=W_{\mu}{}^{\varrho}{}_{\nu}+\tfrac{1}{2}\,(T_{\mu}{}^{\varrho}{}_{\nu}(W)-T_{\mu\nu}{}^{\varrho}(W)+T^{\varrho}{}_{\mu\nu}(W)).$ (A.18)
@@ Appendix B Computation of a BRST cohomology
In order to show this, consider the decomposition $\partial_{\varrho}\,\overline{h}_{\mu\nu}$ in irreducible components.

### v3

@@ 8 BRST formulation
In this concluding section brst rules (2.1) defining covariant fracton gauge theory are formulated in an equivalent way, off-shell realising the embedding of covariant fractons in linearised mhs theory at the level of symmetries.
Consider two doublets $(B_{\mu\nu},\tilde{b}_{\mu\nu})$ and $(r_{\mu},\tilde{\xi}_{\mu})$, with the following transformations
@@
requiring that both $\tilde{b}_{\mu\nu}$ and $r_{\mu}$ are commuting with ghost number zero, so that $B_{\mu\nu}$ is anticommuting with ghost number 1-1 and $\tilde{\xi}_{\mu}$ is anticommuting with ghost number one. Then, using the doublet theorem Brandt:1989rd ; Piguet:1995er , the local $s$-cohomology on the space $\{h^{(\text{fr})}_{\mu\nu},\lambda\}$ and derivatives is equivalent to the local $s$-cohomology on the enlarged space including the two doublets $\{h^{(\text{fr})}_{\mu\nu},\lambda,B_{\mu\nu},\tilde{b}_{\mu\nu},r_{\mu},\tilde{\xi}_{\mu}\}$. If we redefine $\tilde{b}_{\mu\nu}$ and $\tilde{\xi}_{\mu}$ in this way
@@
$s\,b_{\mu\nu}=\partial_{\mu}\,\mathring{\xi}_{\nu}-\partial_{\nu}\,\mathring{\xi}_{\mu},\quad s\,\mathring{\xi}_{\mu}=0.$ (8.3)
@@
The local $s$-cohomology on $\{h_{\mu\nu}^{\text{(fr)}},\lambda\}$ is equivalent to that on $\{B_{\mu\nu},h_{\mu\nu},b_{\mu\nu},r_{\mu},\mathring{\xi}_{\mu},\lambda\}$. In other words, the doublet $(r_{\mu},\tilde{\xi}_{\mu})$ implements the condition $\mathring{\xi}_{\mu}=\tfrac{1}{2}\,\partial_{\mu}\,\lambda$”, which is implied by $r_{\mu}=0$”; and the doublet $(B_{\mu\nu},\tilde{b}_{\mu\nu})$ implements the condition $b_{\mu\nu}=\partial_{\mu}\,r_{\nu}-\partial_{\nu}\,r_{\mu}$”, which is implied by $B_{\mu\nu}=0$”. In this sense, the covariant fracton gauge transformations can be thought as the restriction of linearised diffeomorphisms to the longitudinal ones, as they were firstly introduced Blasi:2022mbl , and covariant fracton gauge theory can be thought as the restriction of the linearised mhs theory in the case in which the antisymmetric part of the linearised vielbein is $\mathrm{d}$-exact. As a consequence, the linearised torsion is cohomologically equivalent to the fracton field strength in this formulation. Indeed, the linearised vielbein reads
@@
This is the reason why the constants $\alpha_{1}$ and $\alpha_{2}$ appear only in the combination $2\,\alpha_{1}+\alpha_{2}$ in embedding covariant fracton gauge theory in linearised mhs theory.
@@ Acknowledgments
The author is grateful to F. Fecit for encouragements and discussions, and he is deeply indebted with C. Imbimbo, who read the first draft of this work, pointing out a mistake and suggesting how improve it, stimulating the investigation in Sections 6 and 7.
@@ Appendix A Weitzenböck geometry within Cartan formalism
In the formalism introduced by Cartan, one considers the vielbein one-form $e^{a}=e_{\mu}{}^{a}\,\mathrm{d}x^{\mu}$ and the spin connection one-form $\omega^{a}{}_{b}=\omega_{\mu}{}^{a}{}_{b}\,\mathrm{d}x^{\mu}$, and defines the torsion two-form $T^{a}=\frac{1}{2}\,T_{\mu}{}^{a}{}_{\nu}\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}$ and the curvature two-form $R^{a}{}_{b}=\frac{1}{2}\,R_{\mu\nu}{}^{a}{}_{b}\,\mathrm{d}x^{\mu}\,\mathrm{d}x^{\nu}$ in the following way:
@@
[…]
$\omega_{\mu}{}^{a}{}_{b}$ serves as a frame connection (for flat indices). The basis or affine connection $\Gamma_{\mu}{}^{\varrho}{}_{\nu}$ (for curved indices) is related to $\omega_{\mu}{}^{a}{}_{b}$ by the requirement that the full covariant derivative of the vielbein, both with respect to $\omega_{\mu}{}^{a}{}_{b}$ and with respect to $\Gamma_{\mu}{}^{\varrho}{}_{\nu}$ (“second vielbein postulate”), vanishes [^16] This is equivalent to define the covariant derivative with respect to the affine connection $\nabla_{\mu}(\Gamma)$ in terms of the covariant derivative with respect to the spin connection $\nabla_{\mu}(\omega)$ in this way: $\nabla_{\mu}(\Gamma)\,(e^{\varrho}{}_{a}\,V^{a})=e^{\varrho}{}_{a}\,\nabla_{\mu}(\omega)\,V^{a}$, for any $V^{a}$. Indeed on one side $\nabla_{\mu}(\Gamma)\,(e^{\varrho}{}_{a}\,V^{a})=\partial_{\mu}\,V^{\varrho}+\Gamma_{\mu}{}^{\varrho}{}_{\nu}\,V^{\nu}$; on the other $e^{\varrho}{}_{a}\,\nabla_{\mu}(\omega)\,V^{a}=\partial_{\mu}\,V^{\varrho}+e^{\varrho}{}_{a}\,(\partial_{\mu}\,e_{\nu}{}^{a}+\omega_{\mu}{}^{a}{}_{b}\,e_{\nu}{}^{b})\,V^{\nu}$, so that we obtain (A.2). VanNieuwenhuizen:1981ae :
@@
[…]
@@
[…]
@@
[…]
@@
A useful tensor is the contorsion, which is defined as the difference between an arbitrary affine connection and the Levi-Civita connection [^18] Using (A.3), we can also write the contorsion in terms of an arbitrary spin connection and the torsionless one: $K_{\mu}{}^{\varrho}{}_{\nu}=(\omega_{\mu}{}^{a}{}_{b}-{\omega^{\textsc{(gr)}}}_{\mu}{}^{a}{}_{b})\,e^{\varrho}{}_{a}\,e_{\nu}{}^{b}$.
$K_{\mu}{}^{\varrho}{}_{\nu}(\Gamma):=\Gamma_{\mu}{}^{\varrho}{}_{\nu}-{\Gamma^{(\textsc{gr})}}_{\mu}{}^{\varrho}{}_{\nu}$, (A.13)
@@
We can relate the Levi-Civita connection, the Weitzenböck connection and the Weitzenböck torsion in the following way. On one side, using (A.13),
$K_{\mu}{}^{\varrho}{}_{\nu}(W)=W_{\mu}{}^{\varrho}{}_{\nu}-{\Gamma^{\textsc{(gr)}}}_{\mu}{}^{\varrho}{}_{\nu}$. (A.16)
@@
$).-K_{\mu}{}^{\varrho}{}_{\nu}(W)=\tfrac{1}{2}\,(T_{\mu}{}^{\varrho}{}_{\nu}(W)-T_{\mu\nu}{}^{\varrho}(W)+T^{\varrho}{}_{\mu\nu}(W$)). (A.17)
@@
$).{\Gamma^{\textsc{(gr)}}}_{\mu}{}^{\varrho}{}_{\nu}=W_{\mu}{}^{\varrho}{}_{\nu}+\tfrac{1}{2}\,(T_{\mu}{}^{\varrho}{}_{\nu}(W)-T_{\mu\nu}{}^{\varrho}(W)+T^{\varrho}{}_{\mu\nu}(W$)). (A.18)
@@ Appendix B Computation of a BRST cohomology
In order to show this, consider the decomposition $\partial_{\varrho}\,\overline{h}_{\mu\nu}$ in irreducible components: [^19] See note 2 for the notation of Young tableaux.
@@ Appendix C Comparison with other conventions
Appendix D Polarisation matrices
See for example VanNieuwenhuizen:1981ae . Choosing the generator of $z$-axis rotation as
$J_{z}=\begin{pmatrix}0&0&0&0\\ 0&0&-i&0\\ 0&i&0&0\\ 0&0&0&0\end{pmatrix}$, (D.1)
the polarisation vectors with helicities $\pm 1$ are
$\varepsilon_{\mu}^{\pm}=\tfrac{1}{\sqrt{2}}\,(0,1,\pm i,0),$ (D.2)
such that
$J_{z}\,\varepsilon_{\mu}^{\pm}=\pm\,\varepsilon_{\mu}^{\pm}.$ (D.3)
Define the polarisation matrices as
$\ignorespaces\lambda_{\mu\nu}^{0}=\varepsilon_{(\mu}^{+}\,\varepsilon_{\nu)}^{-}$ $=\begin{pmatrix}0&0&0&0\\ 0&0&-i&0\\ 0&i&0&0\\ 0&0&0&0\end{pmatrix}$, (D.4a)
$\lambda_{\mu\nu}^{\pm 1}=\tfrac{1}{\sqrt{2}}\,p_{(\mu}\,\varepsilon_{\nu)}^{\pm}$ $=\frac{1}{2}\begin{pmatrix}0&-1&\mp i&0\\ -1&0&0&1\\ \mp i&0&0&\pm i\\ 0&1&\pm i&0\end{pmatrix}$, (D.4b)
$\lambda_{\mu\nu}^{\pm 2}=\varepsilon_{(\mu}^{\pm}\,\varepsilon_{\nu)}^{\pm}$ $=\frac{1}{2}\begin{pmatrix}0&0&0&0\\ 0&1&\pm i&0\\ 0&\pm i&-1&0\\ 0&0&0&0\end{pmatrix}$. (D.4c)
They are such that
$J_{z},\lambda_{\mu\nu}^{s}]=s\,\lambda_{\mu\nu}^{s}.$ (D.5)
Two useful formulas are the following:
$\begin{pmatrix}0&0&0&0\\ 0&a&b&0\\ 0&b&-a&0\\ 0&0&0&0\end{pmatrix}=(a-i\,b)\,\lambda^{+2}+(a+i\,b)\,\lambda^{-2},\quad\begin{pmatrix}0&-a&-b&0\\ -a&0&0&a\\ -b&0&0&b\\ 0&a&b&0\end{pmatrix}=(a-i\,b)\,\lambda^{+1}+(a+i\,b)\,\lambda^{-1}.$ (D.6)